OpenStudy (anonymous):

Solve the compound inequality: -12 <= -2x -4 <2 Answer in interval notation.

OpenStudy (anonymous):

-12 <= -2x -4 <2 i add 4 on all sides: -12+4<= -2x-4+4<2+4 -8<= -2x < 6 then i multiplied all sides by -1/2 4>= x > -3 or -3 <x <=4

OpenStudy (anonymous):

So what would that be in interval notation form?

OpenStudy (anonymous):

-3 <x <=4 is interval

OpenStudy (anonymous):

so the values of x are bigger than -3 or smaller or equal to 4

OpenStudy (anonymous):

Ok.Thanks.

OpenStudy (anonymous):

you're welcome